OpenStudy (anonymous):

Suppose you turn off the engine of your car when the temperature of the engine reaches 190 °F. If the outside temperature is a constant 65 °F, then the temperature T of the engine t minutes after you turn off the engine satisfies the equation below. ln T − 65 125 = −0.07t (a) Solve the equation for T. (b) Find the temperature of the engine 22 minutes after you turn it off. Give your answer to the nearest tenth of a degree.

OpenStudy (anonymous):

\[\ln((T-65)/125)=-0.07t\]

OpenStudy (anonymous):

rise e on both sides\[e^{\ln\frac{(T-65)}{125}}=e^{-0.07t} \]